OpenStudy (erinkb99):

JUST CHECK MY ANSWER:

OpenStudy (erinkb99):

A radar gun was used to record the speed of a runner (in meters per second) during selected times in the first 2 seconds of a race. Use a trapezoidal sum with 4 intervals to estimate the distance the runner covered during those 2 seconds. Give a 2 decimal place answer and include units. t 0 0.5 1.2 1.5 2 v(t) 0 4.5 7.8 8.3 9.0

OpenStudy (erinkb99):

My answer: (b-a)/n (b-a)/n = (2-0)/4 = 0.5 0.5/2=0.25 0+(2*4.5)+(2*7.8)+(2*8.3)+9.0 0+9+15.6+16.6+9 50.2 50.2(0.25)= 12.55

OpenStudy (erinkb99):

@Evoker @zepdrix

OpenStudy (erinkb99):

@mathstudent55 ?

OpenStudy (evoker):

Oh sorry thought someone had responded

OpenStudy (evoker):

Give me a minute to check that form out again.

OpenStudy (mathstudent55):